LOG-IN
Displaying reports 1-20 of 3390.Go to page 1 2 3 4 5 6 7 8 9 10 End
BIGFOOT (Characterization)
Print this report.
MarcEisenmann - 11:10, Friday 02 May 2025 (4015)Get code to link to this report
Comment to Aztec_60_40_1 (Click here to view original report: 4011)

There was some issue in filename saving that was solved.

I also remove second from the filename to avoid too long names.

This morning I found vi stuck at error message while saving the 3rd map.

The input laser current was about 1.7A.

I used some time to check for optimized laser current.

One issue we have is that somehow the powermeter monitoring the input laser power from the reflection of a PBS is not always sensitive to power fluctuations.

This is likely due to too low power making ambient light at the same order of magnitude of laser light.

It seems that from 2A, the input laser power has similar variation as the transmitted laser power over 20mn (Pnorm = 24.38+/-0.06 and 24.35+/-0.06)

This corresponds to about 5mW transmitted and 210uW for monitoring the input laser power.

R&D (FilterCavity)
Print this report.
MichaelPage - 01:40, Friday 02 May 2025 (4013)Get code to link to this report
Misalignment of OPO

During the previous investigation the ppol became very misaligned to the OPO. I checked today and it is still very misaligned.

First I checked some other stuff. SHG and IRMC are ok.

I tried tapping the ppol EOM, which had no effect on the spectrum. Tapping the OPO case causes a little shift but the OPO case seems to be held tightly. Tapping the OPO input mirror holder causes a large shift with a long transient. Considering that the ppol misalignment was a sudden effect I suspect the PZT may have misaligned the input mirror.

I set up measurement of s-pol BAB in transmission of the OPO (p-pol and s-pol are separated by a Faraday isolator PBS just after the OPO dichroic mirror). I used the stick sensor to measure the power of BAB with the power meter setting 8.8 mW range, BW HI (fixed range and BW HI are necessary for sending the power meter signal to the oscilloscope). The BAB and ppol have a very similar transmission spectrum (figure, yellow is ppol, blue is BAB). Since I didn't touch any of the BAB/ppol alignment optics, this means that the s/p-pol IR beams are nominally co-aligned while the OPO is somehow misaligned. I highly doubt it has anything to do with the crystal so as I suspected it seems the input mirror is misaligned. Normally aligning this inpur mirror is a bit of a pain since the OPO is technically a 3 mirror cavity (first you have to align the laser along the crystal axis and then align the input mirror to optimize the mode structure), however, the input mirror is the only thing that has shifted on this path and it looks like the BAB and ppol are still co-aligned, so I think I can realign the input mirror just by trying to recover ppol/BAB mode matching.

Images attached to this report
4013_20250501183825_img1356.jpg
R&D (FilterCavity)
Print this report.
MichaelPage - 01:26, Friday 02 May 2025 (4004)Get code to link to this report
Turning squeezer back on after long break

The work in this elog was done when I got back from Australia but I forgot to post it.

---

In the last update I fixed the stability of the green injection system. The GRMC error signal demodulation signal from DDS2 DAC2 was found to be drifting in phase with respect to other DDS channels. Switching to the free channel DDS2 DAC0 allowed for the GRMC to stay locked for 1 hr+.

The squeezer has been off since I went to Australia so I turned it back on in the hope that nothing is broken. If everything goes well I hoped to check that we can inject high green power to show that the stable operation really is recovered.

PLLs

I turned on the PLLs and loaded the settings from the PLL software. I did not see any PLL sidebands on the CC spectrum analyzer, and while I could see sidebands on the ppol spectrum analyzer, it would not lock. So I measured the signal levels going into the PLL servo.

CC LO in - 21 MHz 8.22 dBm ok
CC Beat in - ~ -1 dBm ok
ppol LO in - 39 MHz 7.9 dBm ok (39 MHz is the current optimal ppol frequency for maximum nonlinear gain)
ppol Beat in - no signal. I could see a ppol beat signal on the ppol PLL monitor cable though.

It turned out that for the CC the spectrum analyzer was just being weird for some reason, and when I changed the frequency settings a few times it showed the PLL sidebands. For ppol, the power supply to the fiber PD that sends the beat signal to the servo was unplugged (not intentionally as part of the scheduled recent electrical shutdown - rather, the connector is quite loose). The signal ppol Beat in becomes - 6 dBm (ok). I loaded the most recent PLL settings in the computer without any changes (20241115PLLCC.txt and same date for ppol, fig 1 and 2, CC is ADF4001) and both PLLs could lock (monitor channel references fig 3 and 4).

Green generation

The SHG became extremely misaligned (fig 5) since I last looked at it before going to Australia (fig 6, mode matching ~ 89.2%). The SHG normally maintains alignment for a very long time. From the adjustment of the steering mirrors the misalignment was in both pitch and yaw. I aligned to (2060 + 294 + 120 mV) ~ 83.3% (fig 7), I couldn't get the other two peaks down any further. It seems the SHG internal alignment could be an issue eventually but for now it's manageable. Generated green is 300 mW before the green FI and 280 mW before the unused 78 MHz EOM.

The GRMC/MZ assembly was also misaligned, but this is normal since the MZ PZT mirror has never been very stable. I realigned to (1290 + 72 + 54 + 18.4 mV) ~ 89.9% (fig 8). As usual it is currently set for 25 mW stable green to OPO which reads 312 mV on the GRMC TRANSMIS OUT monitor channel. PDH is fine (fig 9).

OPO

I sent ppol to the OPO. The alignment looks nominally ok (fig 10) so I didn't measure the mode matching. ppol mode matching was never particularly critical anyway. However, GRMC suddenly unlocked and was a bit out of phase again. I had thought I fixed this problem by changing the DDS channel but it seems there might still be some issues. I checked the DDS2 DAC0 GRMC DEMOD spectrum to see if it was dirty but it seems fine. I reoptimized the GRMC demod phase from 0 to 335 degrees.

With green going to OPO I applied modulation to green phase shifter to check the CC error signal, which directly indicates the amount of nonlinear gain. As shown in figure 11, when GRPS is modulated the CC1 error signal is sinusoidal and the amplitude is close to when I last looked at squeezing (indicated by the cursors on the yellow trace. Blue is CC2). So it seems the nonlinear gain optimization conditions for ppol PLL frequency and OPO temperature haven't changed much.

A short time after I observed the CC error signals I saw that they had disappeared. The OPO transmission spectrum for ppol had suddenly become very misaligned (fig 12), without touching anything on the table. Before realigning anything I checked the laser current/temperature (ok) and then tried tapping optics along the ppol path to see if anything was loose (no loose optics at first check). Then I tried to check BAB to OPO using power meter in transmission of OPO with green blocked. The s-pol transmission spectrum seems to have some low background of junk modes, and then I can see narrow BAB peaks going up and down slowly as if the BAB is being amplitude modulated somewhere. It was getting late and I couldn't determine whether or not the BAB was really misaligned to the OPO.

During this process it seems the IRMC keeps good alignment so the main laser and 88 MHz EOM should be fine.

Images attached to this report
4004_20250501181412_img1329.jpg 4004_20250501181444_img1330copy.jpg 4004_20250501181611_img1335.jpg 4004_20250501181618_img1339.jpg 4004_20250501181653_img1340.jpg 4004_20250501182038_img1162.jpg 4004_20250501182254_img1341.jpg 4004_20250501182451_img1342.jpg 4004_20250501182458_img1343.jpg 4004_20250501182544_img1344.jpg 4004_20250501182549_img1346.jpg 4004_20250501182638_img1350.jpg
BIGFOOT (Cavity)
Print this report.
ClementJacquet - 18:42, Thursday 01 May 2025 (4012)Get code to link to this report
EOM alignment

Here is the report of the installation and alignment of the EOM in the cristallin cavity experiment. The EOM has been installed after the a telescope in order to adapt the beam diameter to the aperture of the EOM. The measurement and the characterization of the beam parameter befor the first lens of the telscope can be found in this logbook : 

The telescope has been design using jamMt the optical parameters are : 3999

f1 = 200mm

f2 = -100mm

d = 18cm

The expected beam radius after the telecope form the simulation of jamMt is presented on the figure 1. in reality, it was necessary to place the second lens a little bit further away, maybe because our estimation of the optical beam parameter berfore the first lens was not really precise.

We measured the profile at the output of this telescope to check if the beam radius was effectively under 1mm of diameter. The result of this measurement is presented on the figure 2. The 0 position is the one of the first lens of the telescope.

The output surface of the cristal is expected to be wedged and separate the beam into two distinct polarization states. we effectively observed two beam far away from the EOM, we added an half weavelate and a quarter waveplate between the two lens for the telescope in order to fully control the polarization states injected inside the EOM. With this method, we were able to recover a single beam i n transmission

Then we tried to align the EOM and reach the maximum power in transmission. the alignment was performed with 16.4mW at the entrance of the EOM but for now we could not get more than 11.3mW in transmission corresponding to 68.9% of the total injected power. Further work is maybe needed to obtain higher power in transmission.

We also observed that the crystal of the EOM was very dirty and maybe a damage. the figure 3 show the crystal inside the EOM. The replacement of this EOM by another (Maybe at another frequency) is currently under discussion.

Images attached to this report
4012_20250501114154_figure1.png 4012_20250501114204_figure2.png 4012_20250501114214_figure3.jpg
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 22:02, Wednesday 30 April 2025 (4011)Get code to link to this report
Aztec_60_40_1

Marc, Shalika

After solving issues of VI, I restarted the laser current. The power at polcam is 2mW now. 

I am starting a map, just to see if everything is fine. The step size of polarizer is 15deg and map count is 6.

To do:

1. The polarizer alignment should be tuned. The light after it is not linear (most of the time). 

2. I want to add a header in the measurement file, to save data about filtering order. 

Comments related to this report
MarcEisenmann - 11:10, Friday 02 May 2025 (4015)

There was some issue in filename saving that was solved.

I also remove second from the filename to avoid too long names.

This morning I found vi stuck at error message while saving the 3rd map.

The input laser current was about 1.7A.

I used some time to check for optimized laser current.

One issue we have is that somehow the powermeter monitoring the input laser power from the reflection of a PBS is not always sensitive to power fluctuations.

This is likely due to too low power making ambient light at the same order of magnitude of laser light.

It seems that from 2A, the input laser power has similar variation as the transmitted laser power over 20mn (Pnorm = 24.38+/-0.06 and 24.35+/-0.06)

This corresponds to about 5mW transmitted and 210uW for monitoring the input laser power.

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 21:41, Wednesday 30 April 2025 (4010)Get code to link to this report
Comment to fully automated vi preparation (Click here to view original report: 3980)

Today while starting measurements, all of a sudden the plotting function stopped working and we were not able to plot anything. At one point I wanted to set all values on front panel as defaullt so I used the option from labview in edit, and I ran the VI. By mistake I set the value of rotator to 560. The VI crashed. Since then the plotting stopped.

The issue was that when we use set all to default, labview also sets non user controlled stuff to default, like loop configurations. The plotting loop had some time contsant and labview overwrote it.  I tried to make the function again with period 100. Now, the loop works fine. This was the only issue for now, but maybe we will have similar issues in future, coz idk what other values might have changed (for other vi). 

I have saved pdf of front end backend as MapBirefringenceVI (in labview scripts), so that we can use for reference in case something weird happens.

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 19:51, Monday 28 April 2025 (4009)Get code to link to this report
Comment to fully automated vi preparation (Click here to view original report: 3980)

The issue of the map with sample has been resolved. In general, when Vis which take instrument handle, like camera, power meter are inside a loop, should use shift register. The handles should be connected using shift register, otherwise we don't pass the handles to the next call of the VI. 

Everything works fine now. We can take entire maps with different polarization, without need to touch anything in between measurements.  Will just optimize speed here and there.

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 19:38, Monday 28 April 2025 (4008)Get code to link to this report
Comment to fully automated vi preparation (Click here to view original report: 3980)

Also, the VI won't work if you don't open the Projecr first. This is because we are using several instance of .Net

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 19:17, Monday 28 April 2025 (4007)Get code to link to this report
Comment to fully automated vi preparation (Click here to view original report: 3980)

We were having several moments of labview crashes. I saw that we had some patch available from NI, so labview has been updated today. 

Other design updates

1. For polarizer rotator it's critical for polarizer to be able to access its NET properly. (I don't know why) If there are too many nested VIs the net have a hard time having the proper memory referencing. I added the polarizer subvi directly on the main  vi. They work fine now. Also, there were several copies of the cage rotator DLL which was causing issues with proper referencing. This can be changed later if it's too annoying. But, I feel it's better to have polarizer controller accessible on main VI, as we might require to tune the initial (first) polarization manually. 

2. The sequence is polarizer rotates, map without sample, map with sample. 

3. Set initial position of map, and polarizer. Such as start making map with initial polarizatiion as 's'.

the only issue that remains is when the first map (without sample) is made everything is fine. When VI moves to make map with sample, all value goes to nan. Then for 2nd map, everything for map without sample is fine. So, there is a repeatable issue, when making map without sample and with sample. 

BIGFOOT (PockelsCells)
Print this report.
ShalikaSingh - 10:22, Tuesday 22 April 2025 (4003)Get code to link to this report
Comment to circuit with transformer (Click here to view original report: 4002)

After looking at the availability of the cores in Japan, and switching to iron powder cores. I came up with the modified design (attached).

The impedance matching is relatively well, from previous design. I don't want to go above 1kV for now. I don't want arcs flying. 

Images attached to this comment
4003_20250422032025_screenshot20250422102110.png
BIGFOOT (PockelsCells)
Print this report.
ShalikaSingh - 09:22, Monday 21 April 2025 (4002)Get code to link to this report
circuit with transformer

I optimized the design of the RF circuit to have better inductor and use transformer to do impedance matching. The core material can be 43. The fig shows more detail about the plan. 

For imepdance matching, the resonant part impedance is 44kohm. To better take into account this value, the Q of the circuit should be first measured.

Since, sqrt(44E3/200) = 15. Therefore, the input transfomer should have turns ratio 1:15. 

Reference materials:

To decide the core material you can use:   https://toroids.info/

and to know how to make RF transformer, use: https://youtu.be/41q0eVpNN8k?si=fb9kXypXaEAJZAcn

Images attached to this report
4002_20250421021911_screenshot20250421091925.png
Comments related to this report
ShalikaSingh - 10:22, Tuesday 22 April 2025 (4003)

After looking at the availability of the cores in Japan, and switching to iron powder cores. I came up with the modified design (attached).

The impedance matching is relatively well, from previous design. I don't want to go above 1kV for now. I don't want arcs flying. 

Images attached to this comment
4003_20250422032025_screenshot20250422102110.png
KAGRA MIR (Spectroscopy)
Print this report.
MarcEisenmann - 00:18, Saturday 19 April 2025 (4001)Get code to link to this report
long term background measurement

[Marc, Shalika, Tsuzuki]

We set up a long term measurement over the week-end without sample (but its holder present) and p-polarized light.

We'll take 1 spectrum / hour.

We also took temperature and humidity sensor from TAMA FC (TSP01) that will provide temperature of the light source, integrating sphere together with temperature and humidity from the table were is located the spectrophotometer.

They will be measured every 900s.

Tsuzuki-san already performed such measurement last year with unpolarized light that showed strong correlation between background transmission and temperature.

Images attached to this report
4001_20250418171802_probe.jpeg 4001_20250418171810_temperature.jpeg
BIGFOOT (Cavity)
Print this report.
MarcEisenmann - 00:12, Saturday 19 April 2025 (4000)Get code to link to this report
Comment to Laser characterisation (Click here to view original report: 3999)

Just to add that the first HWP is slightly tilted so that the reflected beam (~1mm size from sensor card) is at the side of the laser aperture.

The beam sampler was actually replaced by a 2 inch polarizing beam splitter from CVI (TFP-1064-PW-2037).

The previous issue with power control was solved by tuning it closer to its Brewster angle (56deg) allowing to change the transmitted power from 16 to 370mW (maximum power at 2A).

We reduced the laser current to 2A to prevent potential damage of this 30years old laser following Akutsu-san recommendation.

BIGFOOT (Cavity)
Print this report.
ClementJacquet - 16:51, Friday 18 April 2025 (3999)Get code to link to this report
Laser characterisation

Here is the report of the laser characterization for the cavity experiment (effect of birefringence on alignment strategy of crystalline cavity)

-Laser nominal specification :

reference : LightWave 126-1064-500

serial number : 204

Manufactured date : July 1994

Maximum power : 2W

 

-Control specification: 

name : LightWave 125/126

serial number : L-7-2001-6

 

First the relation between the electric intensity provided by the controler and the optical power at the output of the laser has been characterized. The data and the linear fit made with a least square method can be found on the figure 1. The measurent has been done at 26cm from the laser aperture without optical layout between both.

Then we installed a half wavelength plate and a beam sampler in order to controle the power that will be injected in the future optical system. The first optics are know installed, the actual setup can be seen on the figure 2

From this setup, we measured the beam radius as a function of the postion from the lens. The lens is used to focus the beam as the beam radius was too large for the beam profiler aperture, its focal lens is 100mm. Simulation work will be necessary to recover the real beam radius from the laser aperture and will presented in a future logbook or as a reply of this one. The beam radius as a function of the distance from the lens can be seen on the figure 3. The measurement has been done for two radius in two orthogonal directions. The parameter found are :

For Ww(d)

w0 = 263.72 ± 32.54 um

z0 = 11.51 ± 0.08 cm

zr = 1.29 ± 0.18 cm 

 

For Wv(d)

w0 = 278.19 ± 2.68 um

z0 = 11.86 ± 0.09 cm

zr = 1.03 ± 0.19 cm

 

With Ww and Wv the beam radius in two orthogonal direction keping the notation used by the beam profiler software. 

Images attached to this report
3999_20250418094040_figure1.png 3999_20250418094048_figure2.png 3999_20250418094054_figure3.png
Comments related to this report
MarcEisenmann - 00:12, Saturday 19 April 2025 (4000)

Just to add that the first HWP is slightly tilted so that the reflected beam (~1mm size from sensor card) is at the side of the laser aperture.

The beam sampler was actually replaced by a 2 inch polarizing beam splitter from CVI (TFP-1064-PW-2037).

The previous issue with power control was solved by tuning it closer to its Brewster angle (56deg) allowing to change the transmitted power from 16 to 370mW (maximum power at 2A).

We reduced the laser current to 2A to prevent potential damage of this 30years old laser following Akutsu-san recommendation.

BIGFOOT (PockelsCells)
Print this report.
ShalikaSingh - 21:49, Thursday 17 April 2025 (3997)Get code to link to this report
circuit issues

I changed the LC tank circuit, inductor to 470uH. These are the old inductors in elec shop. I couldn't understand why my resonant frequency shifted from 2.1Mhz to 1.63MHz. First I took it as some drift in pockel cell capacitance. But, then I tried to find the specs of the inductor. (It's bad!!)

The self resonance frequency of the inductor is 2.4Mhz. It's literally near my resonant freq. I was not expecting it to be this low. If we use such inductor near their self resonant freq, all forms of parasitic capacitance will affect my circuit. This explains also why my circuit could only achieve half of the desined modulation. The radial leaded inductor stored in elec shop are not good for RF applications. Being so close to self resonance freq can lead to degraded quality factor as well.

I found specs of some from coilcraft, and their self resonance freq is 210Mhz. Some inductor like this will solve the issue. There are some chip based inductors in elec shop, I will check their specs. 

Fig 1 is the circuit currently. The opamp used is EL2099CT. 

File 2 is datasheet of 470uH from elec shop. 

Images attached to this report
3997_20250417151037_screenshot20250417221300.png
Non-image files attached to this report
BIGFOOT (Cavity)
Print this report.
MarcEisenmann - 10:04, Thursday 17 April 2025 (3996)Get code to link to this report
Comment to laser installation (Click here to view original report: 3993)

by rotating the polarization before a pbs or a tilted beam sampler we should be able to tune the reflected (or transmitted) power.

However the typical change we saw (0.3%) is far lower than expected.

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 00:01, Thursday 17 April 2025 (3995)Get code to link to this report
Comment to laser installation (Click here to view original report: 3993)

Why should the power change when you rotate HWP? The polarization from laser is perfectly linear. So, all you do when you rotate hwp is rotate polarization. The power can be rather controlled with a rotating polarizer. 

BIGFOOT (cell birefringence)
Print this report.
MarcEisenmann - 23:26, Wednesday 16 April 2025 (3994)Get code to link to this report
Comment to measurements of gels(ktk693, 697, 699) (Click here to view original report: 3983)

Another issue is that the separate temperature vi was not on.

hence all temperature are 0 on the measurements file..

BIGFOOT (Cavity)
Print this report.
MarcEisenmann - 23:25, Wednesday 16 April 2025 (3993)Get code to link to this report
laser installation

[Clement, Marc]

Yesterday, Shalika and Clement found a spare laser controller nearby TAMA input laser bench that seemed to work properly.

We tried to use it but the current range is between 1.4A to 2.6A.

Clement made power vs current graph that will be uploaded soon.

While the laser manual specifies a waist 5cm after the laser head aperture, it seems to be more around 15cm.

We started to install a power controller (rotating HWP + tilted beam sampler) but somehow the reflected power is almost independent of HWP angle.

We replaced the beam sampler by a TFP polarizing beam splitter and got similar result...

We added a QWP after the HWP and got same result.

Note that we're using (for the first time) a high power power meter head found in PCI (S145).

Comments related to this report
ShalikaSingh - 00:01, Thursday 17 April 2025 (3995)

Why should the power change when you rotate HWP? The polarization from laser is perfectly linear. So, all you do when you rotate hwp is rotate polarization. The power can be rather controlled with a rotating polarizer. 

MarcEisenmann - 10:04, Thursday 17 April 2025 (3996)

by rotating the polarization before a pbs or a tilted beam sampler we should be able to tune the reflected (or transmitted) power.

However the typical change we saw (0.3%) is far lower than expected.

KAGRA MIR (Spectroscopy)
Print this report.
MarcEisenmann - 23:15, Wednesday 16 April 2025 (3992)Get code to link to this report
Comment to polarized measurement (Click here to view original report: 3991)

Today I tried to measure the sample only with p-pol from calibration.

The input polarization is set by rotating the polarizer (spol = 358deg, ppol = 88deg).

It seems I made a mistake in saving the no sample measurement before the sample...

The strange features seen in sample transmission spectrum with ppol is also visible without sample.

More analysis to follow.

The spectrophotometer is off.